# EvalNet: A Practical Toolchain for Generation and Analysis of Extreme-Scale Interconnects

### v3

@@ Appendix A Appendix
Figure 20. Shortest path plot for Cascade Dragonfly, PolarStar and Spectralfly. The amount of shortest paths of length 1 and 2 decreases, if the network gets bigger, because a single node in a group can only reach 21 nodes from another group with at most 2 jumps. Meanwhile, all 95 nodes in the same group can be reached with 2 gaps. The amount of node pairs that have a shortest path $\geq$ 3 stays almost the same. Spectralfly has the least amount of shortest paths of length 1 for all network sizes. The amount of shortest paths of low length stays the same or gets even better for higher-size networks. The Jellyfish equivalent of the Cascade Dragonfly has only a diameter of 3 or 4 (depending on the network size), while Cascade Dragonfly is a diameter 5 topology. Spectralfly’s Jellyfish equilibrium on the other hand looks similar to its base counterpart. Only for large network sizes, the Jellyfish graph shows no shortest paths of length 5, while the Spectralfly graph still has some.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 21. Multiplicity plot for the Cascade Dragonfly, PolarStar and Spectralfly networks. Most node pairs in the Cascade Dragonfly topology have multiple ($\geq$ 5) shortest paths. The multiplicity count either stays about the same or gets bigger, as the Cascade Dragonfly graph increases its size. The small Spectralfly network has only multiplicity counts $\leq$ 4. For large and medium networks, the Spectralfly topology has slightly more multiplicity count for $n_{min}$, however only by a small margin. For small and medium networks, the equivalent Jellyfish graph of the Cascade Dragonfly has as well the most shortest path multiplicity for $n_{min}\geq$ 5. The large Jellyfish has about the same multiplicity count over all values of $n_{min}$. While the equivalent Jellyfish of Spectralfly drops in multiplicity counts for higher values of $n_{min}$, for medium and large networks it increases.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 22. Shortest path plot for Megafly with different values for g and Dragonfly. We can see, that Megafly is a diameter 3 network as well. Megafly has no shortest path of length 1, since every leaf router is only connected to spine routers, and not to another leaf router. Because two nodes have only shortest path of length $l$ = 2 if they are in the same group, most shortest paths are $l_{min}$ = 3 for Megafly. All topologies have the same amount of shortest paths of length 3 for all network sizes. The analysis of the equivalent Jellyfishes looks pretty similar compared to the base Megafly network, where Megafly has shortest paths of length $l$. The Jellyfish graphs have additional shortest paths of length 1 and sometimes length 4.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 23. Multiplicity plot for the Dragonfly and Megafly topologies with different values for g. All topologies have mostly only one shortest path between any two nodes. Between any two leaf nodes in Megafly, there is mostly only one shortest path. There is either one shortest path between the nodes or a bunch ($\geq$ 5) of them. Dragonfly has some shortest paths of different multiplicity, but most lie in the extremes $n_{min}$ = 1 or $n_{min}\geq$ 5. Their equivalent Jellyfish on the other hand have well-distributed shortest path multiplicities.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 24. Shortest path plot for Fat Trees and the k-ary 3-tree. All topologies exhibit the same patterns for the shortest paths of a given length. While the shortest paths of length 4 stay the same for any network size, the number of paths of length 2 decreases as the network gets bigger.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 25. Multiplicity plot for Fat Trees and the k-ary 3-tree. All networks and all sizes have a path multiplicity count of 5 or higher.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 26. Shortest path plot for length $l_{min}$ up to $l$ = 5 for multiple variants of xGFTs. There are only shortest paths of even length. There exist more shortest paths of length 4 than length 2 in general. As the network size increases, the number of shortest paths at each length decreases.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 27. Multiplicity plot for multiple variants of xGFTs. Variant 4 has the lowest amount of multiplicity count. It only has a multiplicity count of 2 and 4. For large network sizes, the 8S variant has even more multiplicity than the smaller networks of that variant.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 28. Edge disjoint path plots for Kautz and Arrangement graphs of small sizes. The connectivity for $A_{n,4}$ is better than the connectivity for $A_{n,5}$. Both have in general fewer edge disjoint paths than the Kautz topologies. $K_{6,n}$ has the best connectivity for these topologies. Furthermore, the Kautz networks analysis looks similar to its equivalent Jellyfish.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 29. Edge disjoint path plots for meshes of different dimensions. For all lengths $l$, the meshes have mostly zero edge disjoint paths between any two nodes. Furthermore, we can see that there are almost no node pairs, where there are 2 paths of a given length $l$. In contrast, none of the Jellyfish equivalent graphs have zero edge disjoint paths for $l$ = 5. The 4D mesh has a slightly worse path diversity than the 5D and 6D meshes.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 30. Edge disjoint path plots for the Tofu topology, 6D torus and 6D Express Mesh. Similar to the diversity analysis of the meshes, all the topologies have zero path diversity for most node pairs of any given length. For both Tofu and the 6D torus network, their path diversity counts lie in three groups. They have diversity counts of 1, $\approx$ 4 and $\approx$ 10. This is even more extreme for the 6D torus network and cannot be seen for the 6D mesh graph. Same as before, the equivalent Jellyfish graphs have a well-distributed diversity count.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 31. Edge disjoint path plots for Cascade Dragonfly, PolarStar and Spectralfly topologies. Both Spectralfly and Cascade Dragonfly have mostly a low diversity count. Cascade Dragonfly has an additional peak with node pairs of a higher diversity count. Even for length $l$ = 5, Spectralfly has only node pairs with 4 edge disjoint paths.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 32. Edge disjoint path plots for Dragonfly, PolarStar and Slim Fly topologies. Slim Fly has the most path diversity by far. It has a diversity count $\geq$ 10 for most nodes, even for paths of length 3. The diversity analysis of PolarStar is pretty similar to Dragonfly for length 1 to 4. For $l$ = 5 we can see that Dragonfly has a higher diversity count by 2 for all node pairs. The PolarStar graph with Paley subgraphs has slightly higher edge disjoint paths for $l$ = 5. Both variants of PolarStar have the same or better diversity compared to their Jellyfish equivalent for all lengths.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 33. Edge disjoint path plots for Dragonfly and Megafly networks with different values of g. All Megafly networks have a low path diversity of either 0 or 1 for most node pairs and all lengths. Furthermore, all have a second peak for a very high diversity count. Depending on the value of g, this peak is between 7 and 15. Although Megafly has node pairs with a higher diversity count, Dragonfly has a well-distributed diversity count between 1 and 5. It is difficult to compare the values with this plot. The diversity counts of all Jellyfish equilibriums show a Gauss distribution for $l$ = 3.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 34. Edge disjoint path plots for the Fat Trees and k-ary 3-tree. All the topologies have only specific values for the path diversity. None of the topologies have any node pairs that have a diversity count of zero for $l\geq$ 4. The k-ary 3-tree has the best path diversity among all other topologies.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 35. Edge disjoint path plots for the xGFTs of different variants. There exist only edge disjoint paths of specific values. For $l\geq$ 4, xGFT8 does not have any edge pairs with a diversity count of zero anymore. The k-ary n-tree has edge nodes with a higher diversity count than any other topology, but it has still node pairs with a diversity count of zero for $l$ = 5.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 36. Low connectivity plot for $K_{4,3}$ Kautz graph of length $l$ = 3. All router pairs have a connectivity $\leq$ 75% for $l$ = 3. We can see the segments (blue) of length 4 connecting a node to its neighbours. This can be seen horizontally as well as vertically. Due to this mesh of edges, all nodes are connected and there exist only a couple of node pairs with connectivity $<$ 25%.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 37. Low connectivity plot for $K_{4,3}$ Kautz graph of length $l$ = 5. All router pairs have a connectivity $\geq$ 75% for $l$ = 5.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 38. Low connectivity plot for the PolarStar graph of length $l$ = 3. The structure graph is an ER2 graph and the subgraph is a BDF3 graph. We can see the 7 well-connected BDF3 clusters of each 8 nodes in the diagonal. The other well-connected areas exist due to the connection of the ER2 graph. Every node in 0 - 7 is connected to a node in 31 - 39, 40 - 47 and 48 - 55.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 39. Low connectivity plot for the PolarStar graph for the length $l$ = 5. The structure graph is an ER2 graph and the subgraph is a BDF3 graph. We can clearly see, that for $l$ = 5 this PolarStar graph has a good connectivity $\geq$ 75% for all router pairs.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 40. Low connectivity plot for the Cascade Dragonfly topology with two clusters for the length $l$ = 3. Each cluster has a good connectivity within itself. On the diagonal, we can also see the 16 chassis for the 5 levels. The connectivity between the two clusters is very low and only about half of the nodes in each cluster have a path to nodes in the other cluster.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 41. Low connectivity plot for the Cascade Dragonfly topology with two clusters for the length $l$ = 5. Right underneath the first cluster, we can see the four node pairs that connect the two clusters. Even for paths of $l$ = 5, the connectivity between the two clusters stays low at only 10%.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 42. Low connectivity plot for the Tofu topology with X = Y = Z = 2 and $l$ = 3. Not all nodes are connected with each other at length $l$ = 3. The 8 clusters of 12 nodes each are visible in the diagonal. The small blue diagonal lines show the edges between the nodes in different clusters. Each row is intersected by 3 of these lines. This is the number of connections each node has to other clusters.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 43. Low connectivity plot for the Tofu topology with X = Y = Z = 2 and length $l$ = 6. There is a good connectivity $\geq$ 70% for most node pairs.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 44. Interference plots for Kautz and Arrangement graphs of length $l$ = 5. The Arrangement networks have the same or less interference than their Jellyfish equivalent counterpart for $l\leq$ 5. For $l$ = 5, both have slightly more interference. The same can be seen for $K_{4,n}$ and $K_{5,n}$ Kautz graphs, but not for $K_{6,n}$, which has the most interference for $l$ = 3 and $l$ = 4, but the least interference for $l$ = 5.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 45. Two-dimensional detailed interference plots for Kautz and Arrangement graphs of length $l$ up to 5. $A_{n,4}$ has more interference than $A_{n,5}$. The different values of k for $K_{b,n}$ Kautz graphs have less of an impact on the amount of interference, but at which length $l$ they will occur. Most of these networks have the most interference at $l$ = 4. Their equivalent Jellyfish have similar interference analysis, but in general a bit less.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 46. Interference plots for meshes of different dimensions of lengths up to $l$ = 5. While 4D and 6D meshes have a pretty similar interference analysis, the 5D mesh has more interference. Furthermore, for $l\leq$ 4, the base networks of all the meshes have higher or equal interference than their equivalent Jellyfish. 4D and 6D meshes have better interference for $l$ = 5, which is not true for the 5D mesh. For $l$ = 5, all meshes have more node pairs with zero interference compared to their equivalent Jellyfish graphs.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 47. Two-dimensional detailed interference plots for meshes of different dimensions of lengths up to $l$ = 5. We can see, that all meshes have similar interferences analysis. For the base mesh networks and $l$ = 5, all components are affected by interference, while for the equivalent Jellyfish, only a subset of them are.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 48. Interference plots for Tofu, 6D mesh and 6D torus with lengths up to $l$ = 5. The Tofu topology has about the same amount of interference as the 6D torus network, but more than the regular 6D mesh. For $l$ = 5, both torus and the Tofu graph have more interference than the 6D mesh and more than any of the Jellyfish equivalents.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 49. Interference plots for PolarStar topology with both subgraphs, Dragonfly and Slim Fly of length $l$ = 1 up to $l$ = 5. Slim Fly has the most interference at $l$ = 3, but less interference for all other values of $l$. The other topologies have the most interference for $l$ = 4. The interference analysis looks similar for the topologies when comparing them to their equivalent Jellyfish, although all of the base topologies have slightly more interference.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 50. Two-dimensional detailed interference plots for PolarStar topology with both subgraphs, Dragonfly and Slim Fly of length $l\leq$ 5. The detailed analysis of both PolarStar graphs are fairly similar. While for Dragonfly, the interference for $l$ = 4 is contributed by a lot of components, only a small amount of components contribute to the high amount of interference for PolarStar.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 51. Interference plots for Cascade Dragonfly, Spectralfly and PolarStar with both subgraphs of length $l$ = 5. Spectralfly has little interference compared to the other topologies for all lengths. It has even about the same amount of interference as its equivalent Jellyfish topology. In contrast to all the other topologies, Cascade Dragonfly has two groups of node quadruples: the ones with little interference and the other group with high interference. Due to the second group of nodes with high interference, Cascade Dragonfly has the most interference of all the topologies. The most extreme difference when comparing to its equivalent Jellyfish, can be seen for $l$ = 4 and $l$ = 5. While the Jellyfish graph has at most interference 3 for any node quadruple, there are groups of nodes with interference values $>$ 15.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 52. Interference plots for Megafly with different gap sizes and Dragonfly of length $\leq$ 5. In general, Megafly has less interference than Dragonfly for length $l\leq$ 3. Only MF8 has a few outlying node pairs with a high interference at $l$ = 3. All Megafly variants have most node pairs with interference between 0 and 3. For $l\geq$ 4, all variants start to have a few $\approx$ 10 outliers with higher interference. This cannot be seen in the Dragonfly topology.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 53. Two-dimensional detailed interference plots for Megafly with different gap sizes and Dragonfly. For lengths of 2 and 3, only a few components contribute to the total interference in the Megafly networks. This changes for $l>$ 3, where almost all values for $c_{l}(\{a,c\},\{b\})+c_{l}(\{a,c\},\{d\})$ are possible. The same cannot be said for the analysis of the Dragonfly graph. For all lengths only a subset of possible values for the components contribute the total interference. This is also true for all Jellyfish equivalent graphs.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 54. Interference plots for k-ary n-trees for different values for k, Fat Trees and Fat Trees2x. For all base topologies, most node pairs have an interference of zero. Both Fat Tree variants have a few outliers that have a higher interference for length 2 and 3. The k-ary 3-tree has these outliers also, although only for $l\geq$ 4. There are also more outlying node pairs than in the Fat Tree variants.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 55. Two-dimensional detailed interference plots for the k-ary 3-tree, Fat Tree and Fat Tree2x. For both components, there exist only specific values they can get. Since values in the diagonal cancel each other out, we can see that there is only one peak next to interference zero for all base topologies.
[…]
[Main findings are in the caption.]Main findings are in the caption.
@@
Figure 56. Interference plots for xGFTs of different variants and k-ary n-tree. Similar to before, all topologies have interference of zero for most node pairs. We can also see the peaks at a specific interference value. xGFT8S has the lowest interference of all of them. It has only interference higher than zero for $l>$ 3 and the peaks have only interference of 3. xGFT8 has only an interference $>$ 0 at length 2 and 3.
[…]
[Main findings are in the caption.]Main findings are in the caption.

### v4

@@ Appendix A Appendix
Figure 20: Shortest path plot for Cascade Dragonfly, PolarStar and Spectralfly. The amount of shortest paths of length 1 and 2 decreases, if the network gets bigger, because a single node in a group can only reach 21 nodes from another group with at most 2 jumps. Meanwhile, all 95 nodes in the same group can be reached with 2 gaps. The amount of node pairs that have a shortest path $\geq$ 3 stays almost the same. Spectralfly has the least amount of shortest paths of length 1 for all network sizes. The amount of shortest paths of low length stays the same or gets even better for higher-size networks. The Jellyfish equivalent of the Cascade Dragonfly has only a diameter of 3 or 4 (depending on the network size), while Cascade Dragonfly is a diameter 5 topology. Spectralfly’s Jellyfish equilibrium on the other hand looks similar to its base counterpart. Only for large network sizes, the Jellyfish graph shows no shortest paths of length 5, while the Spectralfly graph still has some.
@@
Figure 21: Multiplicity plot for the Cascade Dragonfly, PolarStar and Spectralfly networks. Most node pairs in the Cascade Dragonfly topology have multiple ($\geq$ 5) shortest paths. The multiplicity count either stays about the same or gets bigger, as the Cascade Dragonfly graph increases its size. The small Spectralfly network has only multiplicity counts $\leq$ 4. For large and medium networks, the Spectralfly topology has slightly more multiplicity count for $n_{min}$, however only by a small margin. For small and medium networks, the equivalent Jellyfish graph of the Cascade Dragonfly has as well the most shortest path multiplicity for $n_{min}\geq$ 5. The large Jellyfish has about the same multiplicity count over all values of $n_{min}$. While the equivalent Jellyfish of Spectralfly drops in multiplicity counts for higher values of $n_{min}$, for medium and large networks it increases.
@@
Figure 22: Shortest path plot for Megafly with different values for g and Dragonfly. We can see, that Megafly is a diameter 3 network as well. Megafly has no shortest path of length 1, since every leaf router is only connected to spine routers, and not to another leaf router. Because two nodes have only shortest path of length $l$ = 2 if they are in the same group, most shortest paths are $l_{min}$ = 3 for Megafly. All topologies have the same amount of shortest paths of length 3 for all network sizes. The analysis of the equivalent Jellyfishes looks pretty similar compared to the base Megafly network, where Megafly has shortest paths of length $l$. The Jellyfish graphs have additional shortest paths of length 1 and sometimes length 4.
@@
Figure 23: Multiplicity plot for the Dragonfly and Megafly topologies with different values for g. All topologies have mostly only one shortest path between any two nodes. Between any two leaf nodes in Megafly, there is mostly only one shortest path. There is either one shortest path between the nodes or a bunch ($\geq$ 5) of them. Dragonfly has some shortest paths of different multiplicity, but most lie in the extremes $n_{min}$ = 1 or $n_{min}\geq$ 5. Their equivalent Jellyfish on the other hand have well-distributed shortest path multiplicities.
@@
Figure 24: Shortest path plot for Fat Trees and the k-ary 3-tree. All topologies exhibit the same patterns for the shortest paths of a given length. While the shortest paths of length 4 stay the same for any network size, the number of paths of length 2 decreases as the network gets bigger.
@@
Figure 25: Multiplicity plot for Fat Trees and the k-ary 3-tree. All networks and all sizes have a path multiplicity count of 5 or higher.
@@
Figure 26: Shortest path plot for length $l_{min}$ up to $l$ = 5 for multiple variants of xGFTs. There are only shortest paths of even length. There exist more shortest paths of length 4 than length 2 in general. As the network size increases, the number of shortest paths at each length decreases.
@@
Figure 27: Multiplicity plot for multiple variants of xGFTs. Variant 4 has the lowest amount of multiplicity count. It only has a multiplicity count of 2 and 4. For large network sizes, the 8S variant has even more multiplicity than the smaller networks of that variant.
@@
Figure 28: Edge disjoint path plots for Kautz and Arrangement graphs of small sizes. The connectivity for $A_{n,4}$ is better than the connectivity for $A_{n,5}$. Both have in general fewer edge disjoint paths than the Kautz topologies. $K_{6,n}$ has the best connectivity for these topologies. Furthermore, the Kautz networks analysis looks similar to its equivalent Jellyfish.
@@
Figure 29: Edge disjoint path plots for meshes of different dimensions. For all lengths $l$, the meshes have mostly zero edge disjoint paths between any two nodes. Furthermore, we can see that there are almost no node pairs, where there are 2 paths of a given length $l$. In contrast, none of the Jellyfish equivalent graphs have zero edge disjoint paths for $l$ = 5. The 4D mesh has a slightly worse path diversity than the 5D and 6D meshes.
@@
Figure 30: Edge disjoint path plots for the Tofu topology, 6D torus and 6D Express Mesh. Similar to the diversity analysis of the meshes, all the topologies have zero path diversity for most node pairs of any given length. For both Tofu and the 6D torus network, their path diversity counts lie in three groups. They have diversity counts of 1, $\approx$ 4 and $\approx$ 10. This is even more extreme for the 6D torus network and cannot be seen for the 6D mesh graph. Same as before, the equivalent Jellyfish graphs have a well-distributed diversity count.
@@
Figure 31: Edge disjoint path plots for Cascade Dragonfly, PolarStar and Spectralfly topologies. Both Spectralfly and Cascade Dragonfly have mostly a low diversity count. Cascade Dragonfly has an additional peak with node pairs of a higher diversity count. Even for length $l$ = 5, Spectralfly has only node pairs with 4 edge disjoint paths.
@@
Figure 32: Edge disjoint path plots for Dragonfly, PolarStar and Slim Fly topologies. Slim Fly has the most path diversity by far. It has a diversity count $\geq$ 10 for most nodes, even for paths of length 3. The diversity analysis of PolarStar is pretty similar to Dragonfly for length 1 to 4. For $l$ = 5 we can see that Dragonfly has a higher diversity count by 2 for all node pairs. The PolarStar graph with Paley subgraphs has slightly higher edge disjoint paths for $l$ = 5. Both variants of PolarStar have the same or better diversity compared to their Jellyfish equivalent for all lengths.
@@
Figure 33: Edge disjoint path plots for Dragonfly and Megafly networks with different values of g. All Megafly networks have a low path diversity of either 0 or 1 for most node pairs and all lengths. Furthermore, all have a second peak for a very high diversity count. Depending on the value of g, this peak is between 7 and 15. Although Megafly has node pairs with a higher diversity count, Dragonfly has a well-distributed diversity count between 1 and 5. It is difficult to compare the values with this plot. The diversity counts of all Jellyfish equilibriums show a Gauss distribution for $l$ = 3.
@@
Figure 34: Edge disjoint path plots for the Fat Trees and k-ary 3-tree. All the topologies have only specific values for the path diversity. None of the topologies have any node pairs that have a diversity count of zero for $l\geq$ 4. The k-ary 3-tree has the best path diversity among all other topologies.
@@
Figure 35: Edge disjoint path plots for the xGFTs of different variants. There exist only edge disjoint paths of specific values. For $l\geq$ 4, xGFT8 does not have any edge pairs with a diversity count of zero anymore. The k-ary n-tree has edge nodes with a higher diversity count than any other topology, but it has still node pairs with a diversity count of zero for $l$ = 5.
@@
Figure 36: Low connectivity plot for $K_{4,3}$ Kautz graph of length $l$ = 3. All router pairs have a connectivity $\leq$ 75% for $l$ = 3. We can see the segments (blue) of length 4 connecting a node to its neighbours. This can be seen horizontally as well as vertically. Due to this mesh of edges, all nodes are connected and there exist only a couple of node pairs with connectivity << 25%.
@@
Figure 37: Low connectivity plot for $K_{4,3}$ Kautz graph of length $l$ = 5. All router pairs have a connectivity $\geq$ 75% for $l$ = 5.
@@
Figure 38: Low connectivity plot for the PolarStar graph of length $l$ = 3. The structure graph is an ER2 graph and the subgraph is a BDF3 graph. We can see the 7 well-connected BDF3 clusters of each 8 nodes in the diagonal. The other well-connected areas exist due to the connection of the ER2 graph. Every node in 0 - 7 is connected to a node in 31 - 39, 40 - 47 and 48 - 55.
@@
Figure 39: Low connectivity plot for the PolarStar graph for the length $l$ = 5. The structure graph is an ER2 graph and the subgraph is a BDF3 graph. We can clearly see, that for $l$ = 5 this PolarStar graph has a good connectivity $\geq$ 75% for all router pairs.
@@
Figure 40: Low connectivity plot for the Cascade Dragonfly topology with two clusters for the length $l$ = 3. Each cluster has a good connectivity within itself. On the diagonal, we can also see the 16 chassis for the 5 levels. The connectivity between the two clusters is very low and only about half of the nodes in each cluster have a path to nodes in the other cluster.
@@
Figure 41: Low connectivity plot for the Cascade Dragonfly topology with two clusters for the length $l$ = 5. Right underneath the first cluster, we can see the four node pairs that connect the two clusters. Even for paths of $l$ = 5, the connectivity between the two clusters stays low at only 10%.
@@
Figure 42: Low connectivity plot for the Tofu topology with X = Y = Z = 2 and $l$ = 3. Not all nodes are connected with each other at length $l$ = 3. The 8 clusters of 12 nodes each are visible in the diagonal. The small blue diagonal lines show the edges between the nodes in different clusters. Each row is intersected by 3 of these lines. This is the number of connections each node has to other clusters.
@@
Figure 43: Low connectivity plot for the Tofu topology with X = Y = Z = 2 and length $l$ = 6. There is a good connectivity $\geq$ 70% for most node pairs.
@@
Figure 44: Interference plots for Kautz and Arrangement graphs of length $l$ = 5. The Arrangement networks have the same or less interference than their Jellyfish equivalent counterpart for $l\leq$ 5. For $l$ = 5, both have slightly more interference. The same can be seen for $K_{4,n}$ and $K_{5,n}$ Kautz graphs, but not for $K_{6,n}$, which has the most interference for $l$ = 3 and $l$ = 4, but the least interference for $l$ = 5.
@@
Figure 45: Two-dimensional detailed interference plots for Kautz and Arrangement graphs of length $l$ up to 5. $A_{n,4}$ has more interference than $A_{n,5}$. The different values of k for $K_{b,n}$ Kautz graphs have less of an impact on the amount of interference, but at which length $l$ they will occur. Most of these networks have the most interference at $l$ = 4. Their equivalent Jellyfish have similar interference analysis, but in general a bit less.
@@
Figure 46: Interference plots for meshes of different dimensions of lengths up to $l$ = 5. While 4D and 6D meshes have a pretty similar interference analysis, the 5D mesh has more interference. Furthermore, for $l\leq$ 4, the base networks of all the meshes have higher or equal interference than their equivalent Jellyfish. 4D and 6D meshes have better interference for $l$ = 5, which is not true for the 5D mesh. For $l$ = 5, all meshes have more node pairs with zero interference compared to their equivalent Jellyfish graphs.
@@
Figure 47: Two-dimensional detailed interference plots for meshes of different dimensions of lengths up to $l$ = 5. We can see, that all meshes have similar interferences analysis. For the base mesh networks and $l$ = 5, all components are affected by interference, while for the equivalent Jellyfish, only a subset of them are.
@@
Figure 48: Interference plots for Tofu, 6D mesh and 6D torus with lengths up to $l$ = 5. The Tofu topology has about the same amount of interference as the 6D torus network, but more than the regular 6D mesh. For $l$ = 5, both torus and the Tofu graph have more interference than the 6D mesh and more than any of the Jellyfish equivalents.
@@
Figure 49: Interference plots for PolarStar topology with both subgraphs, Dragonfly and Slim Fly of length $l$ = 1 up to $l$ = 5. Slim Fly has the most interference at $l$ = 3, but less interference for all other values of $l$. The other topologies have the most interference for $l$ = 4. The interference analysis looks similar for the topologies when comparing them to their equivalent Jellyfish, although all of the base topologies have slightly more interference.
@@
Figure 50: Two-dimensional detailed interference plots for PolarStar topology with both subgraphs, Dragonfly and Slim Fly of length $l\leq$ 5. The detailed analysis of both PolarStar graphs are fairly similar. While for Dragonfly, the interference for $l$ = 4 is contributed by a lot of components, only a small amount of components contribute to the high amount of interference for PolarStar.
@@
Figure 51: Interference plots for Cascade Dragonfly, Spectralfly and PolarStar with both subgraphs of length $l$ = 5. Spectralfly has little interference compared to the other topologies for all lengths. It has even about the same amount of interference as its equivalent Jellyfish topology. In contrast to all the other topologies, Cascade Dragonfly has two groups of node quadruples: the ones with little interference and the other group with high interference. Due to the second group of nodes with high interference, Cascade Dragonfly has the most interference of all the topologies. The most extreme difference when comparing to its equivalent Jellyfish, can be seen for $l$ = 4 and $l$ = 5. While the Jellyfish graph has at most interference 3 for any node quadruple, there are groups of nodes with interference values >> 15.
@@
Figure 52: Interference plots for Megafly with different gap sizes and Dragonfly of length $\leq$ 5. In general, Megafly has less interference than Dragonfly for length $l\leq$ 3. Only MF8 has a few outlying node pairs with a high interference at $l$ = 3. All Megafly variants have most node pairs with interference between 0 and 3. For $l\geq$ 4, all variants start to have a few $\approx$ 10 outliers with higher interference. This cannot be seen in the Dragonfly topology.
@@
Figure 53: Two-dimensional detailed interference plots for Megafly with different gap sizes and Dragonfly. For lengths of 2 and 3, only a few components contribute to the total interference in the Megafly networks. This changes for $l>$ 3, where almost all values for $c_{l}(\{a,c\},\{b\})+c_{l}(\{a,c\},\{d\})$ are possible. The same cannot be said for the analysis of the Dragonfly graph. For all lengths only a subset of possible values for the components contribute the total interference. This is also true for all Jellyfish equivalent graphs.
@@
Figure 54: Interference plots for k-ary n-trees for different values for k, Fat Trees and Fat Trees2x. For all base topologies, most node pairs have an interference of zero. Both Fat Tree variants have a few outliers that have a higher interference for length 2 and 3. The k-ary 3-tree has these outliers also, although only for $l\geq$ 4. There are also more outlying node pairs than in the Fat Tree variants.
@@
Figure 55: Two-dimensional detailed interference plots for the k-ary 3-tree, Fat Tree and Fat Tree2x. For both components, there exist only specific values they can get. Since values in the diagonal cancel each other out, we can see that there is only one peak next to interference zero for all base topologies.
@@
Figure 56: Interference plots for xGFTs of different variants and k-ary n-tree. Similar to before, all topologies have interference of zero for most node pairs. We can also see the peaks at a specific interference value. xGFT8S has the lowest interference of all of them. It has only interference higher than zero for $l>$ 3 and the peaks have only interference of 3. xGFT8 has only an interference >> 0 at length 2 and 3.